Question 27 of 42 Submit The pH of a 0.0066 M solution of Ca(OH)2 is Х +- x 100
Answer #1

Ca(OH)2 - cat + 2.00 0.0066 M 2X0.0066 M [ÕH)= @X0.0066 M pOH = -log [OH) = -log(2X0.0066) por = 1-88 pH + port = 14 pH = 12.